OpenStudy (anonymous):

What is the rate of change from x = 0 to x = pi over 2?

OpenStudy (zzr0ck3r):

of?

OpenStudy (zzr0ck3r):

unless its a constant function I assume you mean average rate?

OpenStudy (anonymous):

those are the answer possibilities

OpenStudy (zzr0ck3r):

the question does not make sense....

OpenStudy (anonymous):

The amplitude is 3 and the period is 1, so the function is given by y = 3 Sin Θ At Θ₁ = 0, y₁ = 3 Sin 0 y₁ = 0 At Θ₂ = π/2, y₂ = 3 Sin π/2 y₂ = 3 m = (y₂ - y₁) / (Θ₂ - Θ₁) Subbing known values, m = (3 - 0) / (π/2 - 0) m = 3 / π/2 m = 3(2/π) m = 6/π m = 1.91

OpenStudy (anonymous):

thank you so much!

OpenStudy (zzr0ck3r):

this is not the rate of change from 0 to 2pi as that question does not make sense for a non constant function.

OpenStudy (anonymous):

Its really easy actually, To find this, you need to find the slope of the interval. The easiest way to do this is to just do a simple slope equation of rise over run. This is (3/pi/2) =3/pi/2 or 1.910 or 6/pi